OpenStudy (anonymous):

Math Analysis Find the indicated term of the arithmetic sequence. t1=5, d=1/2. Find t21

OpenStudy (anonymous):

\[t _{n}=t _{1}+(n-1)d\] so \[t _{21}=5+(21-1)1/2=5+10=15\]

OpenStudy (anonymous):

ok, thx.. how about t1=-5, r=1/4. Find t4 :)